An article published in the Washington Post claims that 45 percent of all Americans have brown eyes. A random sample of n=76 college students students found 28 who had brown eyes.
Consider testing
H0:p=.45
Ha:p≠.45

Answer by Boreal(15235)

28/76=0.368
z=(0.368-0.45)/sqrt(0.45*0.55/76)
=-0.082/0.0571
=-1.43 (rounded)
the probability z<-1.43 =0.0754 (not rounding) and double it for a two way test, so it is 0.1508. Insufficient evidence to reject Ho.
To get the value of the calculator (0.1528), the point estimate above has to be taken to four decimal places and that used in the formula.
-
can also do normal approximation
mean is np=34.2
variance is np(1-p)=18.81
sd is sqrt (V)=4.34
probability get 28 is z< (28-34.2)/4.34 which ix z<1.43